If a+b+c = 8 and ab + bc + ac = 13 , find the value of (a^2) + (b^2) + (c^2)
Answer by aliafzal14(13)

PROBLEM:
If a+b+c = 8 and ab + bc + ac = 13 , find the value of (a^2) + (b^2) + (c^2)
SOLUTION:
(a+b+c)^2=a^2+2ab+2ac+b^2+2bc+c^2
(8)^2=a^2+b^2+c^2 2(13)
64 - 26 = a^2+b^2+c^2
a^2+b^2+c^2 = 38